A and B are Involved in a duel. The rules of the duel are that they are to shoot each other simultaneously. If one or both are hit, then the duel is over. If shots miss, then they repeat the process. Suppose that the results of the shots are independent and that each shot of A will hit B with probability 0.5, and each shot of B will hit A with probability 0.4

1. What is the probability that the duel ends at the first round?

2. Wha is the probability that the duel ends at the 4th round of shots?

3. Given that the duel ends at the 4th round, what is the probability that A is not hit?
Answer #1

0 Given that, probability o each shot of A will hit B with 0.5 (let us call it a) and each shot of B will hit A with probabthe 1.) m =1, So, the brobability th dual ends at the first round is a 1- ( l-a) (1-6) tes - 1-(1-0.5) (1-0-4) D = 1- 0.5 80-= a (1-6) (1-9) 4-4(1-1) 14 I at last in A is not hit, but - Bishit ? And P(y) = [1- (1-2) (1-2)] (1-2) 4-1 (1-6)Ay So, p(x(Y